OpenStudy (anonymous):

Solve Each Equation. 1/x+1/2x=1/6

OpenStudy (mathlegend):

@Mertsj can you help here? This person isn't online... but I want to know if I am solving this right.

OpenStudy (mertsj):

\[\frac{1}{x}+\frac{1}{2x}=\frac{2}{2x}+\frac{1}{2x}=\frac{3}{2x}\]

OpenStudy (mertsj):

So we have the proportion: \[\frac{3}{2x}=\frac{1}{6}\]

OpenStudy (mertsj):

Cross multiply and solve.

OpenStudy (mathlegend):

Wait but... when you multiplied you multiplied by the 2 in the denominator though right?

OpenStudy (mertsj):

\[\frac{1}{x}\times \frac{2}{2}=\frac{2}{2x}\]

OpenStudy (mertsj):

So I replaced 1/x with 2/2x

OpenStudy (mertsj):

Then since the denominators of the two fractions are the same, you can add them by adding the numerators.

OpenStudy (mathlegend):

\[2(\frac{ 1 }{ 2 }x) = x\]

OpenStudy (mathlegend):

Can I show you what I was trying to do?

OpenStudy (mertsj):

That is true.

OpenStudy (mertsj):

yes

OpenStudy (mathlegend):

\[2(\frac{ 1 }{ x } + \frac{ 1 }{ 2 }x = \frac{ 1 }{ 6 })\]

OpenStudy (mathlegend):

\[\frac{ 2 }{ x } + x = \frac{ 1 }{ 3 }\]

OpenStudy (mathlegend):

Right so far?

OpenStudy (mathlegend):

Could this work?

OpenStudy (mertsj):

I am confused about what the original problem is

OpenStudy (mertsj):

Is it this: \[\frac{1}{x}+\frac{1}{2x}=\frac{1}{6}\]

OpenStudy (mertsj):

or \[\frac{1}{x}+\frac{1}{2}x=\frac{1}{6}\]

OpenStudy (mathlegend):

I believe so.

OpenStudy (mathlegend):

Oh, I mean the 2nd one

OpenStudy (mertsj):

So what you have written is correct.

OpenStudy (mertsj):

\[\frac{2}{x}+x=\frac{1}{3}\]

OpenStudy (mathlegend):

\[x(\frac{ 2 }{ x } + x = \frac{ 1 }{ 3 })\]

OpenStudy (mertsj):

yes

OpenStudy (mathlegend):

\[2+x ^{2}=\frac{ 1 }{ 3 }x\]

OpenStudy (mertsj):

yes

OpenStudy (mathlegend):

Now I am stuck...

OpenStudy (mathlegend):

because I have x^2...

OpenStudy (mertsj):

Multiply both sides by 3

OpenStudy (mathlegend):

\[6 + 3x ^{2} = x\]

OpenStudy (mertsj):

yes. Now subtract x from both sides.

OpenStudy (mathlegend):

\[3x ^{2}-x+6\]

OpenStudy (mathlegend):

set equal to 0

OpenStudy (mertsj):

=0

OpenStudy (mertsj):

Now use the quadratic formula because I don't think it will factor.

OpenStudy (mathlegend):

ok

OpenStudy (mathlegend):

\[\frac{ 1\pm {71i} }{ 6 }\]

OpenStudy (mathlegend):

?

OpenStudy (mertsj):

sqrt71

OpenStudy (mathlegend):

\[\frac{ 1\pm \sqrt{-71} }{ 6 }\]

OpenStudy (mathlegend):

but we can't have a negative under the radical because it is imaginary right?

OpenStudy (mertsj):

\[\frac{1\pm \sqrt{71}i}{6}\]

OpenStudy (mathlegend):

Oh, haha... thank you :)

OpenStudy (mertsj):

yw. But it might have been easier to clear all the fractions right at the beginning by multiplying by the common denominator which is 6x

OpenStudy (mertsj):

\[6x \times\frac{1}{x}\times+6x \times \frac{1}{2}x=6\times\frac{1}{6}\]

OpenStudy (mertsj):

\[6+3x^2=x\]

OpenStudy (mertsj):

\[3x^2-x+6=0\]

OpenStudy (mathlegend):

oh I see

OpenStudy (mertsj):

Good.